OpenStudy (anonymous):

5/8+radical 7 (simplify this)

OpenStudy (joannablackwelder):

I think we just need to find common denominators and add the fractions. Do you know how to do that?

OpenStudy (anonymous):

no I am terrible at math

OpenStudy (anonymous):

I came up with 2,4,6,8,10

OpenStudy (joannablackwelder):

Ok. The common denominator between 8 and 1 is 8. So, we need to give qsrt7 a denominator of 8.

OpenStudy (joannablackwelder):

\[\frac{ 5 }{ 8 }+\frac{ 8\sqrt{7} }{ 8 }\]

OpenStudy (joannablackwelder):

Does that make sense so far?

OpenStudy (anonymous):

Honestly no math is not my strong point nor do I like it I am a Marine

OpenStudy (joannablackwelder):

Ok, thanks for your honesty. To add two fractions together, the fractions must have common denominators. Like to add \[\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }\]

OpenStudy (joannablackwelder):

You first must find common denominators so that we are comparing apples to apples.

OpenStudy (joannablackwelder):

To do that, we must use multiples of the denominators. So, \[\frac{ 1 }{ 2 }\times \frac{ 2 }{ 2 }\] \[=\frac{ 2 }{ 4 }\]gives an equal fraction but with a denominator of 4.

OpenStudy (anonymous):

common denominator between these two is 4

OpenStudy (joannablackwelder):

Multiplying by \[\frac{ 2 }{ 2 }\] doesn't change the value, just the way it is written.

OpenStudy (joannablackwelder):

You got it! :)

OpenStudy (joannablackwelder):

Once you find common denominators, add the numerators and leave the denominators the same.

OpenStudy (anonymous):

so 2,4,6,8,10?

OpenStudy (joannablackwelder):

I don't understand what you mean.

OpenStudy (anonymous):

SOOOOOOO sorry we were simplifying the radical?

OpenStudy (joannablackwelder):

For my example, \[\frac{ 2 }{ 4 }+\frac{ 1 }{ 4 }=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

I came up with 40-5sqrt7/57

OpenStudy (joannablackwelder):

How did you get that?

OpenStudy (anonymous):

google

OpenStudy (joannablackwelder):

Let's try to do it by hand. :)

OpenStudy (joannablackwelder):

Does it make more sense now how I got \[\frac{ 5 }{ 8 }+\frac{ 8\sqrt{7} }{ 8 }\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

google was right huh... I have a TBI so me understanding math is not good

OpenStudy (joannablackwelder):

No, I'm not quite sure where that came from.

OpenStudy (joannablackwelder):

\[\sqrt{7}\] can be written as a fraction by \[\frac{ \sqrt{7} }{ 1 }\]

OpenStudy (anonymous):

thats not even one of the choices

OpenStudy (joannablackwelder):

I'm not done yet. :)

OpenStudy (joannablackwelder):

So now we have \[\frac{ 5 }{ 8 }+\frac{ \sqrt{7} }{ 1 }\]

OpenStudy (joannablackwelder):

And to add the fractions we need to find common denominators.

OpenStudy (joannablackwelder):

Making any more sense?

OpenStudy (anonymous):

so then whats the final?

OpenStudy (joannablackwelder):

I hadn't gotten there yet, and when you left, I saw no need to keep going if you weren't following.

OpenStudy (anonymous):

I went and ate dinner

OpenStudy (anonymous):

I'm ready coach

OpenStudy (joannablackwelder):

No worries. :)

OpenStudy (joannablackwelder):

Ok, so let's recap. Do you remember that we are trying to combine the terms because 1 term is considered more simple than 2?

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

Cool. And to combine the terms, we need to add the fractions which means we need common denominators.

OpenStudy (joannablackwelder):

Do you follow me?

OpenStudy (anonymous):

sorry I got kicked

OpenStudy (joannablackwelder):

Kicked?

OpenStudy (anonymous):

out

OpenStudy (joannablackwelder):

haha, no worries.

OpenStudy (joannablackwelder):

Do you understand that in order to combine the terms we need to add them together? And to add them together, we need to add fractions?

OpenStudy (anonymous):

umm yes

OpenStudy (joannablackwelder):

So, is that a kinda?

OpenStudy (anonymous):

how is this all simplifying 5/8+sqrt7

OpenStudy (anonymous):

kinda no I have a brain injury and this crap causes me to get angry

OpenStudy (joannablackwelder):

:( I'll try to make it as simple as possible.

OpenStudy (anonymous):

lol you want simple then tell me the answer ha

OpenStudy (joannablackwelder):

As simple for you to understand as possible....

OpenStudy (joannablackwelder):

:)

OpenStudy (anonymous):

tell me was I right with the answer then?

OpenStudy (joannablackwelder):

No, I don't think so. Not yet.

OpenStudy (joannablackwelder):

\[\frac{ 5 }{ 8 }+\frac{ \sqrt{7} }{ 1 } \] is as far as we got. You can write any number as itself over 1 and that is the same thing.

OpenStudy (joannablackwelder):

To find a common denominator, we need to multiply the second fraction by 8/8 which doesn't change the value but gives us the common denominator.

OpenStudy (joannablackwelder):

\[\frac{ 5 }{ 8 }+\frac{ \sqrt{7} }{ 1 }\times \frac{ 8 }{ 8 }\]

OpenStudy (joannablackwelder):

\[=\frac{ 5 }{ 8 }+\frac{ 8\sqrt{7} }{ 8 }\]

OpenStudy (joannablackwelder):

Making sense?

OpenStudy (anonymous):

now we just take the 5 and 8 and multiply them together?

OpenStudy (joannablackwelder):

No, now that we have common denominators, we add the numerators (the top) and leave the denominator the same.

OpenStudy (joannablackwelder):

\[\frac{ 5+8\sqrt{7} }{ 8 }\]

OpenStudy (joannablackwelder):

And I think that is as far as the simplification can go...

OpenStudy (joannablackwelder):

Make sense?

OpenStudy (anonymous):

A. 5/57 b. 5radical 7/15 c. 5-radical 7/8 d. 40-5sqrt7/57

OpenStudy (joannablackwelder):

Wow, really? Those are pretty crazy!

OpenStudy (joannablackwelder):

You could just punch it all into your calculator and see which one gives you the same decimal.

OpenStudy (anonymous):

huh

OpenStudy (joannablackwelder):

Put 5/8 +sqrt7 in your calculator and get a decimal.

OpenStudy (anonymous):

yea 3.27051311?

OpenStudy (joannablackwelder):

Cool. Now put the answer choices into your calculator and get decimals.

OpenStudy (joannablackwelder):

Then see which one matches. :)

OpenStudy (joannablackwelder):

Looks right to me. :)

OpenStudy (anonymous):

This one is a pain

OpenStudy (anonymous):

I despise word problems

OpenStudy (joannablackwelder):

This is a common ratio problem. tall/long = tall/long It's not so bad. :)

OpenStudy (joannablackwelder):

\[\frac{ 6.9 }{ 20.7 }=\frac{ 15.9 }{ ? }\]

OpenStudy (anonymous):

When you are just trying to get it done then it is a problem

OpenStudy (joannablackwelder):

Yeah, true.

OpenStudy (joannablackwelder):

Do you know how to solve this kind?

OpenStudy (anonymous):

no

OpenStudy (joannablackwelder):

Replace the ? with an x (I like to work with variables better at least).

OpenStudy (anonymous):

I'm just going to make an educated guess lol 29.7 feet

OpenStudy (joannablackwelder):

I think it will need to be bigger than that. Between 3 and 4 times the height.

OpenStudy (anonymous):

ok so 47.7

OpenStudy (joannablackwelder):

That sounds more like it.

OpenStudy (anonymous):

really?

OpenStudy (joannablackwelder):

Yup. It is the same common ratio.

OpenStudy (joannablackwelder):

To get to 20.7, you have to multiply 6.9 by 3.

OpenStudy (joannablackwelder):

So, to find the unknown side, multiply 15.9 by 3 too.

OpenStudy (joannablackwelder):

Good job!

OpenStudy (joannablackwelder):

Oh, is this one the one from before? That is why it is different. I thought the sqrt 7 was a different term, not in the denominator....

OpenStudy (anonymous):

different